# BugMagnifier: TON Transaction Simulator for Revealing Smart Contract Vulnerabilities

### v1

@@ IV-A Choosing a Blockchain Emulator
Our evaluation revealed critical limitations in existing tools. TON Contract Executor [18] abstracts away all phases except compute, bypassing essential TVM opcodes such as ACCEPT_MSG_VALUE and STORING. This creates dangerous blind spots for vulnerabilities involving storage fees or bounce message handling. Similarly, TVM-linker [19] implements storage fees as fixed overhead rather than through TON’s dynamic cell depth calculations and omits bounce phase processing entirely. Both tools fail to preserve intermediate TVM states between phase transitions, rendering them incapable of detecting state inconsistencies that occur during partial execution sequences.
TON Sandbox [20] emerged as the only viable foundation due to its transparent implementation of TON’s transaction lifecycle. Its transaction processor module exposes the complete TVM state machine with precise gas accounting at each opcode execution, preserving full context of intermediate states between phase transitions. Crucially, its modular architecture provides hooks into the message processing pipeline, enabling BugMagnifier’s core innovation: a message interceptor that dynamically reorders messages while maintaining visibility into all five transaction phases. This capability allows systematic exploration of message sequences that trigger temporal vulnerabilities, particularly race conditions where contract behavior depends on the interleaving of messages from multiple senders.
@@ IV-B BugMagnifier
The BugMagnifier interactive console is a powerful tool that provides developers with the ability to emulate TON transactions locally and identify vulnerabilities in smart contracts through advanced message management and simulation of various execution scenarios. The tool is publicly available on GitHub [21]. Built as a TypeScript application leveraging Node.js runtime, BugMagnifier implements a layered architecture with distinct components for contract compilation, transaction simulation, state management, and analysis. At its core, the tool establishes a bidirectional communication channel with the TON Sandbox through a custom IPC (Inter-Process Communication) protocol that enables fine-grained control over transaction execution while maintaining compatibility with the TVM’s deterministic execution requirements.
@@
The console provides comprehensive message queue control through commands in the table I.
TABLE I: BugMagnifier Command Reference
@@ V Illustrative Vulnerable Smart Contract
As part of the illustrative numerical experiment, we employ the BugMagnifier tool to demonstrate the complexities of detecting race conditions in TON smart contracts. To facilitate this analysis, we designed a deliberately vulnerable FunC smart contract implementing a deposit pool mechanism that exhibits critical race conditions under concurrent message processing. This example reveals how subtle implementation details in asynchronous blockchain environments can lead to severe financial inconsistencies, even in seemingly straightforward contracts. The vulnerability stems from improper state management during the ownership assignment phase, a critical weakness that becomes apparent only when analyzing message sequence permutations.
The test contract implements a deposit pool mechanism centered around two core state variables: a 64-bit integer representing the accumulated balance (balance) and a slice storing the owner’s address (owner_address). Upon initial deployment, the contract establishes a zero balance and a null owner address (represented by an empty slice). Transaction processing logic is implemented through two primary operations: OP_ENLIST (value 1) for balance deposits and OP_CLAIM (value 2) for withdrawals. The critical vulnerability emerges from the contract’s ownership assignment mechanism, which triggers during the first deposit operation. This implementation pattern, while seemingly logical for single-user initialization, creates a fundamental race condition when multiple deposit messages arrive concurrently.
The following listing presents the complete vulnerable contract implementation with critical sections annotated for clarity:
@@
7 slice ds = get_data()
8 .begin_parse();
9 int balance = slice_bits(ds) >= 64 ?
10 ds.load_uint(64) : 0;
11 slice owner_address = ds;
@@
13}
14
15(slice) get_msg_sender(cell in_msg) {
16 slice cs = in_msg.begin_parse();
17 int flags = cs.load_uint(4);
18 slice sender_address =
19 cs^.load_msg_addr();
20 return sender_address;
21}
22
23() _save_state(int balance,
24 slice owner_address) impure {
25 set_data(begin_cell()
26 .store_uint(balance, 64)
27 .store_slice(owner_address)
28 .end_cell()
29 );
30}
31
32() recv_internal(int msg_value,
33 cell in_msg, slice in_msg_body)
34 impure {
35 var (balance, owner_address) =
36 _load_state();
37 slice sender_address =
38 get_msg_sender(in_msg);
39
@@
41 int op = in_msg_body
42 .load_uint(32);
@@
45 // CRITICAL VULNERABILITY POINT
46 // Ownership assignment occurs ONLY
47 // during first deposit when
48 // owner_address is empty (0 bits)
@@
50 _save_state(balance,
51 sender_address);
@@
53 _save_state(balance,
54 owner_address);
@@
59 // Ownership verification
60 // before withdrawal
61 if (equal_slices(owner_address,
62 sender_address)) {
@@
64 _save_state(balance, null());
@@
69}
70
71() recv_external(slice _) {
@@
73}
74
75(int, slice) get_state() method_id {
76 return _load_state();
77}
The vulnerability manifests through non-deterministic message ordering during concurrent execution, a direct consequence of TON’s asynchronous architecture. Consider the following scenario involving three messages:
@@
    OP_ENLIST from Alice (sender ID 1)
@@
    OP_ENLIST from Bob (sender ID 2)
@@
    OP_CLAIM from Alice (sender ID 1).
When these messages arrive concurrently, TON’s message queue processing creates two distinct execution paths with divergent outcomes:
Scenario 1: Bob’s message processes first
@@
    Bob’s OP_ENLIST executes first
@@
    Contract detects empty owner_address (0 bits)
@@
    Bob’s address becomes permanent owner
@@
    Alice’s subsequent OP_CLAIM fails ownership verification
  5. 5.
    Final state: balance = 20,000,000 nanotokens, owner = Bob.
Scenario 2: Alice’s message processes first
@@
    Alice’s OP_ENLIST executes first
@@
    Contract detects empty owner_address
@@
    Alice’s address becomes permanent owner
@@
    Alice’s OP_CLAIM succeeds (ownership verified)
  5. 5.
    Final state: balance = 0, owner = null.
This race condition stems from the contract’s dependence on the slice_bits(owner_address) check during ownership assignment, a state transition that should be atomic but is instead exposed to message interleaving. The vulnerability is particularly insidious because:
  • It passes standard functional tests with sequential message processing.
  • It only manifests under specific message order permutations.
  • The financial impact (complete loss of funds) occurs without explicit error conditions.
  • It exploits TON’s asynchronous execution model where message processing order is non-deterministic.
From a technical perspective, the vulnerability arises during the compute phase of transaction execution when the TVM processes the conditional ownership assignment. The critical instruction sequence occurs at the bytecode level as:
1DUP // Duplicate owner_address slice
2SLICELEN // Calculate bit length
3PUSHINT 0 // Compare with zero
4IFEQ // If empty, assign new owner
This sequence creates a race window where concurrent messages can both pass the empty check before either completes state update, a classic check-then-act vulnerability amplified by TON’s message queue architecture.
BugMagnifier enables systematic detection of this vulnerability through its message permutation capabilities. By generating all possible execution sequences of the three-message scenario (6 permutations), the tool identifies two distinct final states:
  1. 1.
    States where Alice becomes owner and withdraws funds (3 permutations)
  2. 2.
    States where Bob becomes owner and funds remain locked (3 permutations).
The differential state analysis reveals critical inconsistencies in both the balance and ownership fields, with the diff command showing:
1State difference detected:
2- balance: 20000000 vs 0
3- owner_address: [Bobs address] vs [empty]
This case study exemplifies how TON’s asynchronous execution model transforms seemingly benign state transitions into critical vulnerabilities. The contract’s failure to implement proper synchronization mechanisms, such as nonce-based ownership claims or explicit state locking, demonstrates the heightened security requirements for asynchronous blockchain development. Our numerical experiments (Section VI) quantify how detection complexity varies with message ratios, revealing that such vulnerabilities become increasingly difficult to identify as message counts grow asymmetrically, a direct consequence of TON’s probabilistic message processing order.
@@ VI Numerical Experiments
The methodology evaluated race condition detection complexity across message ratios with fixed Bob messages ($n_{2}=32$) and geometrically increasing Alice messages ($n_{1}=1,2,4,\dots,512$). For each configuration, For each configuration, 100 independent trials recorded the iteration count where state divergence first occurred, providing a quantitative measure of a contract’s susceptibility to race conditions. The experimental results are presented in Table II.
TABLE II: Experimental Results
@@
The data reveal a U-shaped relationship between the message ratio and the detection complexity. Minimum detection iterations (3.06) occur at balanced message counts ($n_{1}=n_{2}=32$), while extreme imbalances require significantly more iterations (29.32 for $n_{1}=1$, 15.46 for $n_{1}=512$). Higher standard deviations at imbalanced ratios indicate greater variance in the detection timing.
@@
Figure 1 plots experimental results against $\log_{2}(n_{1}/n_{2})$ with error bars representing $\pm$1 standard deviation. The x-axis spans from -5 (Bob dominance) to 4 (Alice dominance), with zero representing balanced messages. The experimental data aligns precisely with the theoretical curve , confirming the probabilistic model of race condition manifestation.

### v2

@@ IV-A Choosing a Blockchain Emulator
Our evaluation revealed critical limitations in existing tools. TON Contract Executor [TONContractExecutor] abstracts away all phases except compute, bypassing essential TVM opcodes such as ACCEPT_MSG_VALUE and STORING. This creates dangerous blind spots for vulnerabilities involving storage fees or bounce message handling. Similarly, TVM-linker [TVMLinker] implements storage fees as fixed overhead rather than through TON’s dynamic cell depth calculations and omits bounce phase processing entirely. Both tools fail to preserve intermediate TVM states between phase transitions, rendering them incapable of detecting state inconsistencies that occur during partial execution sequences.
TON Sandbox [Sandbox] emerged as the only viable foundation due to its transparent implementation of TON’s transaction lifecycle. Its transaction processor module exposes the complete TVM state machine with precise gas accounting at each opcode execution, preserving full context of intermediate states between phase transitions. Crucially, its modular architecture provides hooks into the message processing pipeline, enabling BugMagnifier’s core innovation: a message interceptor that dynamically reorders messages while maintaining visibility into all five transaction phases. This capability allows systematic exploration of message sequences that trigger temporal vulnerabilities, particularly race conditions where contract behavior depends on the interleaving of messages from multiple senders.
@@ IV-B BugMagnifier
The BugMagnifier interactive console is a powerful tool that provides developers with the ability to emulate TON transactions locally and identify vulnerabilities in smart contracts through advanced message management and simulation of various execution scenarios. The tool is publicly available on GitHub [BugMagnifier]. Built as a TypeScript application leveraging Node.js runtime, BugMagnifier implements a layered architecture with distinct components for contract compilation, transaction simulation, state management, and analysis. At its core, the tool establishes a bidirectional communication channel with the TON Sandbox through a custom IPC (Inter-Process Communication) protocol that enables fine-grained control over transaction execution while maintaining compatibility with the TVM’s deterministic execution requirements.
@@
The console provides comprehensive message queue control through commands in the table II.
TABLE II: BugMagnifier Command Reference
@@ V Illustrative Vulnerable Smart Contract
We employ BugMagnifier to demonstrate the detection of race conditions inherent to TON’s asynchronous execution model. For this analysis, we designed a system of smart contracts implementing a deposit pool with a critical vulnerability that manifests under concurrent message processing. This example illustrates how the non-deterministic ordering of messages in TON can lead to severe financial inconsistencies. The vulnerability originates from non-atomic state management during ownership assignment, a flaw that becomes apparent only when examining permutations of messages from independently deployed user agents.
The test system comprises a vulnerable main contract and multiple user agent contracts, modeling real-world TON applications where users interact via their own contract instances, potentially deployed across different shards.
The core of the system is a main contract that manages a deposit pool using two primary state variables: a 64-bit integer ‘balance‘ and a slice owner_address. Upon deployment, the balance is zero and the owner address is empty. The contract processes two operations: OP_ENLIST (value 1) for deposits and OP_CLAIM (value 2) for withdrawals. Each user (e.g., Alice and Bob) operates through a UserContract agent, which stores the main contract’s address and forwards deposit requests.
The vulnerability is in the ownership assignment logic within the main contract’s OP_ENLIST handler. Ownership is assigned to the sender of the first deposit message. This pattern creates a race condition when multiple user agents concurrently send their first deposit messages, as their processing order is non-deterministic in TON’s sharded architecture.
The following listings present the vulnerable contracts.
Main Contract (MainContract.fc):
@@
7 slice ds = get_data().begin_parse();
8 int balance = 0;
9 slice owner_address = "";
10 // Load shared state: balance and owner
11 if (slice_bits(ds) >= 64 + 267) {
12 balance = ds~.load_uint(64);
13 owner_address = ds~.load_msg_addr();
14 }
@@
16}
17
18() recv_internal(int msg_value, cell in_msg, slice in_msg_body) impure {
19 var (balance, owner_address) = _load_state();
20 slice sender_address = get_msg_sender(in_msg);
@@
22 int op = in_msg_body~.load_uint(32);
23 // DEPOSIT OPERATION: balance increment and owner assignment
@@
26 // VULNERABILITY: Check-then-act on shared state ‘owner_address‘.
27 // Two concurrent messages can both pass the empty check before either updates the state.
@@
29 _save_state(balance, sender_address);
@@
31 _save_state(balance, owner_address);
@@
37 if (equal_slices(owner_address, sender_address)) {
38 send_raw_message(begin_cell()
39 .store_uint(0x18, 6)
40 .store_slice(sender_address)
41 .store_coins(balance)
42 .store_uint(0, 1 + 4 + 4 + 64 + 32 + 1 + 1)
43 .store_uint(OP_CLAIM, 32)
44 .end_cell(),
45 64
46 );
@@
48 _save_state(balance, "");
@@
53}
Listing 1: Vulnerable Main Contract with Inter-Contract Race Condition
User Agent Contract (UserContract.fc):
1#include "imports/stdlib.fc";
2const OP_SEND_ENLIST = 1;
3
4() send_enlist(slice main_address, int amount) impure {
5 cell msg = begin_cell()
6 .store_uint(0, 6) // internal, no bounce
7 .store_slice(main_address) // MainContract address
8 .store_coins(amount) // deposit amount
9 .store_uint(0, 1 + 4 + 4 + 64 + 32 + 1 + 1)
10 .store_uint(OP_ENLIST, 32) // operation for MainContract
11 .end_cell();
12 send_raw_message(msg, 3); // mode 3: carry remaining balance
13}
14
15() recv_internal(int msg_value, cell in_msg, slice in_msg_body) impure {
16 if (slice_bits(in_msg_body) >= 32) {
17 int op = in_msg_body~.load_uint(32);
18 if (op == OP_SEND_ENLIST) {
19 int amount = in_msg_body~.load_uint(128);
20 slice main_address = _load_main_address();
21 send_enlist(main_address, amount); // Cross-contract invocation
@@
23 }
24 }
25}
Listing 2: User Agent Contract for Inter-Shard Communication
Consider a scenario with three contract instances:
  • UserContract for Alice (address addr_alice)
  • UserContract for Bob (address addr_bob)
  • MainContract (address addr_main)
Both Alice and Bob simultaneously initiate their first deposits via their agents, generating the following messages to the main contract:
@@
    OP_ENLIST from addr_alice
@@
    OP_ENLIST from addr_bob
@@
    OP_CLAIM from addr_alice (attempting to withdraw the pooled balance)
Due to TON’s asynchronous architecture, the processing order of messages 1 and 2 is non-deterministic, leading to divergent financial outcomes:
Scenario A: Message 1 (Alice) processes first.
@@
    Alice’s deposit executes. owner_address is empty, so Alice becomes the owner.
@@
    Bob’s deposit executes. owner_address is now non-empty, so Bob is not set as owner.
@@
    Alice’s claim executes. Ownership verification succeeds; Alice withdraws the full balance.
@@
    Final state: Balance = 0, Owner = (empty).
Scenario B: Message 2 (Bob) processes first.
@@
    Bob’s deposit executes. owner_address is empty, so Bob becomes the owner.
@@
    Alice’s deposit executes. owner_address is now non-empty; ownership remains with Bob.
@@
    Alice’s claim executes. Ownership verification fails; the withdrawal is rejected.
@@
    Final state: Balance = (total deposits), Owner = addr_bob. Alice loses her funds.
This is a classic check-then-act race condition. The check slice_bits(owner_address) == 0 and the subsequent state update are not atomic. Concurrent messages from different senders can interleave between these steps, causing the contract to assign ownership based on a transient, outdated state view.
The vulnerability is subtle because:
  • It passes linear, single-user functional testing.
  • It manifests only under specific message orderings.
  • It results in a complete, silent loss of funds for one party.
  • It is a direct consequence of TON’s non-deterministic inter-shard message ordering.
BugMagnifier systematically uncovers this flaw through permutation testing. For the three-message sequence, it explores all $3!=6$ possible orderings. The diff analysis reveals two distinct final state classes corresponding to the two scenarios above, clearly identifying the non-deterministic outcome and the critical divergence in the owner_address field. This case study exemplifies how asynchronous execution transforms simple state transitions into critical vulnerabilities, necessitating tools like BugMagnifier for detection.
@@ VI Numerical Experiments
The methodology evaluated race condition detection complexity across message ratios with fixed Bob messages ($n_{2}=32$) and geometrically increasing Alice messages ($n_{1}=1,2,4,\dots,512$). For each configuration, For each configuration, 100 independent trials recorded the iteration count where state divergence first occurred, providing a quantitative measure of a contract’s susceptibility to race conditions. The experimental results are presented in Table III.
TABLE III: Experimental Results
@@
The data reveal a U-shaped relationship between the message ratio and the detection complexity. Minimum detection iterations (3.06) occur at balanced message counts ($n_{1}=n_{2}=32$), while extreme imbalances require significantly more iterations (29.32 for $n_{1}=1$, 15.46 for $n_{1}=512$). Higher standard deviations at imbalanced ratios indicate a greater variance in detection timing.
@@
Figure 1 plots the experimental results against $\log_{2}(n_{1}/n_{2})$ with error bars representing $\pm$1 standard deviation. The x-axis spans from -5 (Bob dominance) to 4 (Alice dominance), with zero representing balanced messages. The experimental data aligns precisely with the theoretical curve , confirming the probabilistic model of race condition manifestation.